The safe stress , for a hollow steel column which carries an axial load of 2200 kN is 120 MN/m^2. If the external diameter of the column is 25cm, determine the internal diameter.



1
Expert's answer
2021-02-01T09:29:57-0500

Given axial load, "P= 2.2\times10^3kN"

Stress "\delta= 120MN/m^2=120N/mm^2=12N/cm"

External diameter = 25cm

Area, A ="\frac{Load}{Stress}=\frac{2.2\times10^3}{12}=183.333"

But the area of hollow A ="\frac{\pi}{4}(D^2-d^2)=183.333\implies \frac{\pi}{4}(25^2-d^2)=183.333"

"0.785714285\times(625-d^2)=183.333\implies d^2=391.6667089"

"d=\sqrt(391.6667089)=19.79 cm"